OpenStudy (anonymous):

√5x-1 = √4x+ 9 solve the equation

OpenStudy (anonymous):

X=10?

OpenStudy (anonymous):

thank you so much could you help with one more

OpenStudy (anonymous):

was that a guess or did you figure it out

OpenStudy (anonymous):

√5x-1 = √4x+ 9 => 5x-1=4x+9

OpenStudy (anonymous):

ok i see wht you are doing? How about this one 2√6 + 3√96

OpenStudy (anonymous):

-1 and 9 under the √??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if not then (√5-√4)*√x=9 √x=9/(√5-√4)=9(√5+√4)

OpenStudy (anonymous):

they are under the √

OpenStudy (anonymous):

could you help me with the other one

OpenStudy (anonymous):

of course wut is it

OpenStudy (anonymous):

2√6 + 3√96

OpenStudy (anonymous):

2√6 + 3√96=2√6+3*4*√6=14√6

OpenStudy (anonymous):

ty so much if your up to it i have a few more

OpenStudy (anonymous):

ok is it assignments at school?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(8+√11) (8-√11)

OpenStudy (anonymous):

4=√p-2

OpenStudy (anonymous):

(8+√11) (8-√11)=8^2-11=5 4=√p-2 =>16=p-2 p=18

OpenStudy (anonymous):

ok that one might not be right because its not a option

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

sorry made a mistake...8^2-11=53

OpenStudy (anonymous):

ok that one is there :-)

OpenStudy (anonymous):

but I dont know whats wrong with the next one what are those options?

OpenStudy (anonymous):

√6, 36, 3, 6

OpenStudy (anonymous):

i have 2 more if you dont mind i will leave you a medal

OpenStudy (anonymous):

if -2 is under √ then 36 is correct

OpenStudy (anonymous):

bring them on!...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lol √3 + 4√3

OpenStudy (anonymous):

what is the domain of the function? y=√3x+3

OpenStudy (anonymous):

5√3 x->0 since the number under √ has to be higher than or equal to 0

OpenStudy (anonymous):

which one is that??

OpenStudy (anonymous):

√3 + 4√3=5√3 the domain x->0

OpenStudy (anonymous):

for that one i just need the equation figured out

OpenStudy (anonymous):

the domain one is y=√3x+3

OpenStudy (anonymous):

thank you for your smarts!! How do i give u a medal??

OpenStudy (anonymous):

thats enough?

OpenStudy (anonymous):

just click best response tk:D

OpenStudy (anonymous):

i still need the domain one if you can

OpenStudy (anonymous):

what is the domain of the function y=√3x+3

OpenStudy (anonymous):

for y=√(3x+3) the domain is x->-3 i dont know whats wrong with the previous answer

OpenStudy (anonymous):

the options are x< -1, x>1, x>-1

OpenStudy (anonymous):

x>-1 then though i think x=-1 is also acceptable

OpenStudy (anonymous):

ok thank you I gave you a medal, you really helped me out

OpenStudy (anonymous):

glad :D